OpenStudy (anonymous):

find all points ob the curve x^2y^2+xy=2 where the slope of the tangent line is -1. Ok so i know you find y prime first, but then what? Thanks in advance...

zepdrix (zepdrix):

Did you try taking a derivative yet? :o What'd you come up with? is it really ugly?

OpenStudy (anonymous):

y prime is: -y-2xy^2/2yx^2+x

zepdrix (zepdrix):

Hmm you're missing a 2 in your numerator I think.

OpenStudy (anonymous):

so the two doesnt go to zero when you take the derivative?

zepdrix (zepdrix):

Oooo good call.. XD

zepdrix (zepdrix):

\[\Large\bf\sf y'\quad=\quad \frac{-y-2xy^2}{2yx^2+x}\]Bahhh ya caught me :D

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

Gotta watch my math there, heh.

OpenStudy (anonymous):

so i know how to find the equation if given the points but i dont know how to find them with only a slope..any ideas?

zepdrix (zepdrix):

Ok so instead of giving a point, they gave a slope. The slope is the value for y'.\[\Large\bf\sf -1\quad=\quad \frac{-y-2xy^2}{2yx^2+x}\]

zepdrix (zepdrix):

And then we have to try and uhhhh.. do something with this.

OpenStudy (anonymous):

ok that seems right...i think lol

zepdrix (zepdrix):

ooo this works out really nicely. I won't spoil the fun. I bet you'll figure it out! :O

OpenStudy (anonymous):

so i assume that you times the denominator by the -1 to get -2yx^2-x=-y-2xy^2, then the 2yx^2 cancel

OpenStudy (anonymous):

x=y+2xy^2

OpenStudy (anonymous):

and y= x-2xy^2

zepdrix (zepdrix):

what? I'm confused. If you cancel the 2xy^2, aren't you just left with -x=-y ?

OpenStudy (anonymous):

oh my gosh yes oops lol

OpenStudy (anonymous):

so then do you plug in the slope to find points?

zepdrix (zepdrix):

so we've found that our curve has slope -1 along y=x. Let's see where our curve `intersects` the line y=x to find the actual points.

zepdrix (zepdrix):

So we'll plug x in for all of our y's.

zepdrix (zepdrix):

\[\Large\bf\sf x^2y^2+xy=2\qquad\quad\to\quad\qquad x^2(x^2)+x(x)=2\]

OpenStudy (anonymous):

to find x's?

OpenStudy (anonymous):

so X^4+X^2=2

zepdrix (zepdrix):

ok good. Understand how to solve for x? It's a little tricky.

OpenStudy (anonymous):

so i assume you bring the 2 over then equal it to 0?

zepdrix (zepdrix):

yes you subtract 2 from each side, leaving zero on the right.

OpenStudy (anonymous):

ok then it looks like X^4+x^2-2=0

OpenStudy (anonymous):

i dont think i know how to factor this one:(

zepdrix (zepdrix):

Let, \(\large\bf\sf u=x^2\)

zepdrix (zepdrix):

\[\Large\bf\sf u^2+u-2=0\]

zepdrix (zepdrix):

can you factor it now? :o

zepdrix (zepdrix):

No. (u-1)(u+2)=0

OpenStudy (anonymous):

but what about the x^4?

zepdrix (zepdrix):

You can't factor out an x^2 from each term. Is that what you were trying to do? Just do the substitution I mentioned.

OpenStudy (anonymous):

so are we not factoring X^4+x^2-2=0? cause there is an x^4

zepdrix (zepdrix):

4th degree polynomials are not easy to factor unless you can immediately understand what is going on. So I was making a substitution to reduce the function to a QUADRATIC. We can factor it easily once it's a quadratic. We will undo the substitution later.

zepdrix (zepdrix):

u=x^2

zepdrix (zepdrix):

u^2 = x^4

zepdrix (zepdrix):

x^4+x^2-2=0 --> u^2+u-2=0

OpenStudy (anonymous):

oh i see what ya did there